OpenStudy (anonymous):

(5b) 2nd power 5x5 = 25 25b 2 to the second power?

OpenStudy (mayankdevnani):

\[\huge (5b)^2=5 \times 5 \times b \times b\]

OpenStudy (mayankdevnani):

can you solve it? @mister88

OpenStudy (anonymous):

25 b to the 2nd power?

OpenStudy (mayankdevnani):

and always remember that the sign of the power is ` ^` Got it.... like `5b^2 `

OpenStudy (mayankdevnani):

understood? @mister88

OpenStudy (anonymous):

I don't understand the question $$(5b)^2=25b^2$$ If thats what your asking, note that the function x^s is multiplicative with respect to the variable x. A fun fact, this property is key to the euler product representation of the reimann zeta function. $$\zeta(s)=\prod_{p}\frac{1}{1-p^{-s}}$$

OpenStudy (anonymous):

Yes, and thanks for the confirmation.

OpenStudy (mayankdevnani):

see @mister88

OpenStudy (mayankdevnani):

understood or not? @mister88

OpenStudy (anonymous):

Yes...

OpenStudy (mayankdevnani):

good